# Flexible Deep Neural Networks for Partially Linear Survival Data

### v1

@@ 8 Appendix - Proofs
Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (1). Because $f(t,{\bf X},{\bf Z})$ is predictable, the stochastic integral $\int_{0}^{\tau}f(t,{{\bf X}},{{\bf Z}})\,dM(t)$ has expectation zero. Consequently,
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
@@
$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{0}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (10)
@@
$0<c_{\lambda}\;\leq\;h_{0}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{\lambda}<\infty.$
Applying these bounds to the right-hand side of (10) yields
$c_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{\lambda}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
@@
Thus (7) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
@@
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (11)
@@
Under Assumptions A1 and A2, the event-time hazard $h_{0}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity. Similarly, Assumption A4 implies the censoring mechanism is well behaved. Therefore,
@@
Using these bounds inside the conditional expectation (11), we obtain, that given $({\bf X},{\bf Z})$,
@@
Thus (8) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
Finally, suppose $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$. Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–a.s. on $\{\Delta=1\}$. From (7) we also have
@@
so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $P_{0}$–almost surely. Likewise, from (8),
@@
$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes P_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
@@
where $\chi_{\eta}$ is the linear predictor and $h_{\eta}(t,{\bf X},{\bf Z})$ is the corresponding instantaneous hazard process. Let $\eta_{0}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{0}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
$\eta_{\varepsilon}=\eta_{0}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d},\tilde{{\bf d}},M)$. Because $\chi_{\eta}$ is linear in $(g,\mbox{${\theta}$})$, the Gateaux derivative of $\chi_{\eta}$ in direction $w$ at $\eta_{0}$ is simply
@@
$h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})\}=h_{o}(t,{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
so at $\varepsilon=0$ we have $h_{\eta_{0}}=h_{o}$ and
$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
@@
$\ell_{0}^{\prime}(\eta_{0})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t,{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
@@
the Doob–Meyer decomposition for the true model (1) gives
$dN(t)=Y(t)h_{o}(t,{\bf X},{\bf Z})\,dt+dM(t)\,$,
@@
$\ell_{0}^{\prime}(\eta_{0})[w]=0\qquad\text{for every direction }w\,.$
In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{0}}(t,{\bf X},{\bf Z})$, we can summarize this step as
$\ell_{0}^{\prime}(\eta_{0})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t,{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
To control $\ell_{0}(\eta)-\ell_{0}(\eta_{0})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$. Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{0}+v(\eta-\eta_{0})$ for $v\in[0,1]$. The mean value theorem yields
$\ell_{0}(\eta)=\ell_{0}(\eta_{0})+\ell_{0}^{\prime}(\eta_{0})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{0}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{0})[w]=0$, it remains to analyse the second derivative term.
@@
which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{0}$ gives
@@
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t,{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
@@
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
@@
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
Next, for relating the second derivative to $d(\eta,\eta_{0})^{2}$, we apply Lemma 1 for
$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\,.$
@@
$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t,{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big\}\,.$
@@
$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t,{\bf X},{\bf Z})}{h_{o}(t,{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
@@
$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T,{\bf X},{\bf Z})}{h_{o}(T,{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
@@
Recall that the distance $d(\eta,\eta_{0})$ is defined by
$d(\eta,\eta_{0})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{0}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
@@
$C_{\star}\,d(\eta,\eta_{0})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{0})^{2}\,.$
@@
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{0})^{2}$. Thus,
$\ell_{0}(\eta)-\ell_{0}(\eta_{0})\;\asymp\;-\,d(\eta,\eta_{0})^{2},$
@@
    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-g^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
    where $g^{\ast}$ denotes the efficient projection defined in Theorem 3.
@@
$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\,\}\,.$
@@
$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(P)}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$. The parametrization $\eta=(\mbox{${\theta}$},g)$ is $s$-sparse in the neural-network coordinates, and Assumption A4 ensures that the weights are uniformly bounded and that the architecture has fixed depth and layer widths. In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in zhong2022deep. Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(P)$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5. Since Lipschitz images of a function class cannot increase its covering number by more than a multiplicative constant, the metric entropy of the localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ is therefore controlled by the entropy of the underlying sparse neural-network ball
$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{0}}:d(\eta,\eta_{0})\leq\delta\}.$
Lemma 6 of zhong2022deep establishes that for such a sparse network class,
@@
where $\mathcal{U}$ is the architectural complexity factor. Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(P)$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union. Consequently,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(P)\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
@@
Finally, applying Lemma 3.4.2 of vdvwellner1996 to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(P)}\,\log N_{[]}(\delta)\Bigr\}\,.$
Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(P)}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality. ∎

### v2

@@ Appendix A Proofs for Section 4
Let $N(t)=\Delta\,\mathbf{1}(T\leq t)$ be the counting process associated with the event time, and let $M(t)=N(t)-\int_{0}^{t}Y(s)\,h_{o}(s\,|\,{{\bf X}},{{\bf Z}})\,ds$ denote its martingale decomposition under the true hazard (2.1). Because $f(t,{\bf X},{\bf Z})$ is predictable, the stochastic integral $\int_{0}^{\tau}f(t,{{\bf X}},{{\bf Z}})\,dM(t)$ has expectation zero. Consequently,
$\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,dN(t)\right\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{o}(t\,|\,{\bf X},{\bf Z})\,dt\right\}.$
@@
$\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}=\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,h_{o}(t\,|\,{\bf X},{\bf Z})\,dt\right\}\,.$ (A.1)
@@
$0<c_{h}\;\leq\;h_{o}(t\,|\,{\bf X},{\bf Z})\;\leq\;C_{h}<\infty.$
Applying these bounds to the right-hand side of (A.1) yields
$c_{h}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\;\leq\;\mathbb{E}\{\Delta\,f(T,{\bf X},{\bf Z})\}\;\leq\;C_{h}\,\mathbb{E}\!\left\{\int_{0}^{\tau}f(t,{\bf X},{\bf Z})\,Y(t)\,dt\right\}\,.$
@@
Thus (4.2) holds with $c_{1}\leq 1$ and $c_{2}=1/\delta_{1}$.
@@
$\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\big\}=\mathbb{E}_{{\bf X},{\bf Z}}\!\Big[\mathbb{E}\!\big\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\,\big|\,{\bf X},{\bf Z}\big\}\Big]\,$, (A.2)
@@
Under Assumptions A1 and A2, the event-time hazard $h_{o}(t,{\bf X},{\bf Z})$ is bounded away from zero and infinity. Similarly, Assumption A4 implies the censoring mechanism is well behaved. Therefore,
@@
Using these bounds inside the conditional expectation (A.2), we obtain, that given $({\bf X},{\bf Z})$,
@@
Thus (4.3) holds with $c_{3}=1/C_{f}$ and $c_{4}=1/c_{f}$ (or equivalently with $c_{3},c_{4}$ chosen directly as above).
Finally, suppose $\mathbb{E}\{\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\}=0$. Since $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}\geq 0$, this implies $\Delta\,\phi(T,{\bf X},{\bf Z})^{2}=0$ almost surely, hence $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–a.s. on $\{\Delta=1\}$. From (4.2) we also have
@@
so $\mathbb{E}\{\phi(T,{\bf X},{\bf Z})^{2}\}=0$, which forces $\phi(T,{\bf X},{\bf Z})=0$ $\mathbb{P}_{0}$–almost surely. Likewise, from (4.3),
@@
$\chi_{\eta_{1}}(t,{\bf X},{\bf Z})=\chi_{\eta_{2}}(t,{\bf X},{\bf Z})\quad\text{for }dt\otimes\mathbb{P}_{{\bf X},{\bf Z}}\text{–almost all }(t,{\bf X},{\bf Z})\,.$
@@
where $\chi_{\eta}$ is the linear predictor and $h_{\eta}(t,{\bf X},{\bf Z})$ is the corresponding instantaneous hazard process. Let $\eta_{o}=(g_{o},\theta_{o})$ denote the true parameter, with true hazard $h_{o}(t,{\bf X},{\bf Z})=h_{\eta_{0}}(t,{\bf X},{\bf Z})$.
We study the behavior of $\ell_{0}$ along a one–dimensional path through $\eta_{o}$. Fix an arbitrary direction $w=(w_{g},w_{\theta})$ and define
$\eta_{\varepsilon}=\eta_{o}+\varepsilon w\,,\qquad\varepsilon\in\mathbb{R}\,.$
Assume that $\varepsilon$ is restricted to a small neighbourhood of 0 so that $\eta_{\varepsilon}\in\mathcal{\eta}\in\Theta\times\mathcal{H}(q,\boldsymbol{\alpha},{\bf d}^{*},\tilde{{\bf d}},M)$. Because $\chi_{\eta}$ is linear in $(g,\mbox{${\theta}$})$, the Gateaux derivative of $\chi_{\eta}$ in direction $w$ at $\eta_{0}$ is simply
@@
$h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})=h_{o}(t\mid{\bf X},{\bf Z})\,\exp\{\chi_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})\}=h_{o}(t\mid{\bf X},{\bf Z})\,\exp\{\varepsilon\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\}\,$,
so at $\varepsilon=0$ we have $h_{\eta_{o}}=h_{o}$ and
$\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,$,
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{\varepsilon}}(t,{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})^{2}\,.$
@@
$\ell_{0}^{\prime}(\eta_{o})[w]$ $=\frac{d}{d\varepsilon}\mathbb{E}\Big\{\Delta\,\chi_{\eta_{\varepsilon}}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\Big|_{\varepsilon=0}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,\frac{d}{d\varepsilon}h_{\eta_{\varepsilon}}(t\mid{\bf X},{\bf Z})\Big|_{\varepsilon=0}\,dt\Big\}$
$=\mathbb{E}\Big\{\Delta\,\dot{\chi}_{w}(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)\,h_{o}(t\mid{\bf X},{\bf Z})\,\dot{\chi}_{w}(t,{\bf X},{\bf Z})\,dt\Big\}\,.$
@@
the Doob–Meyer decomposition for the true model (2.1) gives
$dN(t)=Y(t)h_{o}(t\mid{\bf X},{\bf Z})\,dt+dM(t)\,$,
@@
$\ell_{0}^{\prime}(\eta_{o})[w]=0\qquad\text{for every direction }w\,.$
In particular, when we fix $\eta$ and set $q(t,{\bf X},{\bf Z})=\chi_{\eta}(t,{\bf X},{\bf Z})-\chi_{\eta_{o}}(t,{\bf X},{\bf Z})$, we can summarize this step as
$\ell_{0}^{\prime}(\eta_{o})[w]=\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})-\int_{0}^{\tau}Y(t)h_{o}(t\mid{\bf X},{\bf Z})\,q(t,{\bf X},{\bf Z})\,dt\Big\}=0.$
To control $\ell_{0}(\eta)-\ell_{0}(\eta_{o})$, we apply a one–dimensional Taylor expansion of $\ell_{0}(\eta_{\varepsilon})$ around $\varepsilon=0$. Write $\eta=\eta_{1}$ and define the line segment $\eta_{v}=\eta_{o}+v(\eta-\eta_{o})$ for $v\in[0,1]$. The mean value theorem yields
$\ell_{0}(\eta)=\ell_{0}(\eta_{o})+\ell_{0}^{\prime}(\eta_{o})[w]+\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, where $w=\eta-\eta_{o}$. Since we have already shown that $\ell_{0}^{\prime}(\eta_{o})[w]=0$, it remains to analyse the second derivative term.
@@
which is again linear in $w$. Repeating the differentiation steps above but now starting from $\eta_{v}$ instead of $\eta_{o}$ gives
@@
$\frac{d^{2}}{d\varepsilon^{2}}h_{\eta_{v}+\varepsilon w}(t\mid{\bf X},{\bf Z})\Big|_{\varepsilon=0}=h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,q_{v}(t,{\bf X},{\bf Z})^{2}\,.$
@@
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q_{v}(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\,.$
@@
$\ell_{0}^{\prime\prime}(\eta_{v})[w,w]=-\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\,.$
Next, for relating the second derivative to $d(\eta,\eta_{o})^{2}$, we apply Lemma 1 for
$f_{v}(t,{\bf X},{\bf Z})=q(t,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(t\mid{\bf X},{\bf Z})}{h_{o}(t\mid{\bf X},{\bf Z})}\,.$
@@
$\mathbb{E}\Big\{\int_{0}^{\tau}Y(t)\,q(t,{\bf X},{\bf Z})^{2}\,h_{\eta_{v}}(t\mid{\bf X},{\bf Z})\,dt\Big\}\asymp\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T\mid{\bf X},{\bf Z})}{h_{o}(T\mid{\bf X},{\bf Z})}\Big\}\,.$
@@
$0<c_{\star}\;\leq\;\frac{h_{\eta_{v}}(t\mid{\bf X},{\bf Z})}{h_{o}(t\mid{\bf X},{\bf Z})}\;\leq\;C_{\star}<\infty,\qquad\text{for all }(t,{\bf X},{\bf Z}),\;v\in[0,1]\,.$
@@
$c_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big]\;\leq\;\mathbb{E}\Big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\,\frac{h_{\eta_{v}}(T\mid{\bf X},{\bf Z})}{h_{o}(T\mid{\bf X},{\bf Z})}\Big]\;\leq\;C_{\star}\,\mathbb{E}\big[\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
@@
Recall that the distance $d(\eta,\eta_{o})$ is defined by
$d(\eta,\eta_{o})^{2}=\mathbb{E}\Big\{\Delta\big\{\chi_{\eta}(T,{\bf X},{\bf Z})-\chi_{\eta_{o}}(T,{\bf X},{\bf Z})\big\}^{2}\Big\}=\mathbb{E}\big\{\Delta\,q(T,{\bf X},{\bf Z})^{2}\big\}\,.$
@@
$C_{\star}\,d(\eta,\eta_{o})^{2}\;\leq\;\ell_{0}^{\prime\prime}(\eta_{v})[w,w]\;\leq\;-c_{\star}\,d(\eta,\eta_{o})^{2}\,.$
@@
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})=\frac{1}{2}\,\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$
for some $\bar{v}\in(0,1)$, and we have just shown that $\ell_{0}^{\prime\prime}(\eta_{\bar{v}})[w,w]$ is comparable to $d(\eta,\eta_{o})^{2}$. Thus,
$\ell_{0}(\eta)-\ell_{0}(\eta_{o})\;\asymp\;-\,d(\eta,\eta_{o})^{2},$
@@
    $\psi_{\eta}(T,\mathbf{X},\mathbf{Z},\Delta)=\int_{0}^{T}\bigl\{\mathbf{Z}-{\bf g}^{\ast}(t,\mathbf{X})\bigr\}\,dM_{\eta}(t),$
    where $\bf g$$^{\ast}$ denotes the efficient projection defined in Theorem 3.
@@
$\mathcal{F}_{\delta}=\{\,f_{\eta}-f_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\,\}\,,\qquad\mathcal{G}_{\delta}=\{\,\psi_{\eta}-\psi_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\,\}\,.$
@@
$\|\psi_{\eta_{1}}-\psi_{\eta_{2}}\|_{L^{2}(\mathbb{P})}\;\lesssim\;d(\eta_{1},\eta_{2})\,.$
Hence both localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(\mathbb{P})$-ball of radius $O(\delta)$ and admit a common envelope of order $O(\delta)$. The parametrization $\eta=(\mbox{${\theta}$},g)$ is $s$-sparse in the neural-network coordinates, and Assumption A4 ensures that the weights are uniformly bounded and that the architecture has fixed depth and layer widths. In particular, the map $\eta\mapsto\chi_{\eta}$ belongs to the class of sparse feedforward neural networks considered in Zhong et al. (2022). Moreover, the maps $\eta\longmapsto f_{\eta}$ and $\eta\longmapsto\psi_{\eta}$ are Lipschitz transformations of $\chi_{\eta}$ in the $L^{2}(\mathbb{P})$-metric, with Lipschitz constant bounded uniformly under Assumptions A1A3 and A5. Since Lipschitz images of a function class cannot increase its covering number by more than a multiplicative constant, the metric entropy of the localized classes $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ is therefore controlled by the entropy of the underlying sparse neural-network ball
$\mathcal{H}_{\delta}=\{\chi_{\eta}-\chi_{\eta_{o}}:d(\eta,\eta_{o})\leq\delta\}.$
Lemma 6 of Zhong et al. (2022) establishes that for such a sparse network class,
@@
where $\mathcal{U}$ is the architectural complexity factor. Because $\mathcal{F}_{\delta}$ and $\mathcal{G}_{\delta}$ lie in a common $L^{2}(\mathbb{P})$-ball of radius $O(\delta)$ and share a common envelope of order $O(\delta)$, the same entropy bound applies to their union. Consequently,
$\log N_{[]}\!\left(\varepsilon,\,\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta},\,L^{2}(\mathbb{P})\right)\;\lesssim\;s\log\!\left(\frac{\mathcal{U}\delta}{\varepsilon}\right)\,.$
@@
Finally, applying Lemma 3.4.2 of van der Vaart and Wellner (1996) to the empirical process $\mathbb{G}_{n}=\sqrt{n}(\mathbb{P}_{n}-\mathbb{P})$, with envelope $F_{\delta}=O(\delta)$, gives
$\mathbb{E}^{\ast}\!\left\{\sup_{\mu\in\mathcal{F}_{\delta}\cup\mathcal{G}_{\delta}}|\mathbb{G}_{n}\mu|\right\}\;\leq\;C_{3}\Bigl\{J_{[]}(\delta)+n^{-1/2}\,\|F_{\delta}\|_{L^{2}(\mathbb{P})}\,\log N_{[]}(\delta)\Bigr\}\,.$
Substituting the bound on $J_{[]}(\delta)$ and noting that $\|F_{\delta}\|_{L^{2}(\mathbb{P})}=O(\delta)$ and $\log N_{[]}(\delta)\lesssim s\log(\mathcal{U}/\delta)$ yields the asserted maximal inequality. ∎